Use a proportion to answer the question in terms of y.
What percent of 6y is 24?
Please provide rules for answering questions of this type. My daughter left her book at school, so I can not look at it to determine the way to work it.
Thank You

Answer by stanbon(75887)

What percent of 6y is 24
----------------------------
EQUATION:
(x/100)6y = 24
Proportion:
x/100 = 24/6y
x/100 = 4/y
x = 400/y
==============
(400/y)% of (6y) = 24
=========================
Cheers,
Stan H.